# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ 3. Hypotheses
We study the effect of various biases in a predictive text system on co-writing creative stories. Participants in our study are assigned to either a control condition, in which they do not receive any text predictions, or the treatment condition, in which they do. In the treatment condition, similar to standard phone keyboard interfaces, the participant is provided (up to) three predicted “next words” that they can select rather than typing on their own. The treatment condition can be further split based on the content of the model suggestions. Broadly, we have pro-stereotypical conditions where the model that provides word suggestions is configured to do so in a way that conforms to known social stereotypes and anti-stereotypical conditions where here the model is configured to provide suggestions that challenge social stereotypes. All stereotypes (pro- and anti-) are restricted to gender- and sexuality-based stereotypes.
@@
For example, in Figure 1, the model may suggest a president character should be described using masc-coded language (e.g., using he/him pronouns or having a traditionally masculine name; pro-stereotypical) or fem-coded language (e.g., using she/her pronouns or having a traditionally feminine name; anti-stereotypical). Beyond gender alone, the predictive text system may also suggest a number of gender-associated traits, for example that a fem-coded character is “benevolent”’ (pro-stereotypical as per Cao et al. (2022b)) or that she is “threatening” (anti-stereotypical as per Cao et al. (2022b)).
@@
Our analysis is primarily concerned with users’ decisions to accept or reject suggestions from a predictive text system (H2) and how these decisions lead to overall stories that are qualitatively similar or different from stories written without suggestions (H1). Measures of the acceptance of individual word-level suggestions capture different effects than measures of the degree of the use of stereotypes in the completed. The former provide measures of reliance. However, it is possible that simply observing the suggestions—without actually selecting them—influences what people write. Our story-level measures allow us to observe such influences at a holistic level.
@@
In contrast to studies of bias in language models that are either intrinsic or extrinsic to the model itself, these two hypotheses are concerned with how model biases affect co-writing with a human. H2 focuses on individual micro decisions about when participants accept model suggested words or reject them and write new words, and H1 focuses on the impact of those decisions to written stories more broadly.
@@
Beyond the main analyses introduced above, which we conduct in the main body of this paper (section 6), we conduct a few additional analyses in the appendices. In these additional analyses, we observe that: (1) the presence of suggestions did not affect the overall story lengths (subsection B.2); (2) participants’ gender identity did not significantly affect their acceptance of gendered suggestions (subsection B.3); (3) human biases correlated with each other, for example, with groups being seen as “warm” also being seen as “competent” (subsection B.4); and (4) writing with predictive text did not significantly affect gender gaps in toxicity but led to significant gender gaps in sentiment and character agency in some writing scenarios, with fem-coded characters sometimes being portrayed more positively, yet with less agency. (subsection B.5).
@@ 4. Study Design
The study is conducted using a custom-built mobile web interface (3(a)) mimicking a smartphone keyboard with predictive text (participants were not allowed to use their devices’ built-in keyboard), and participants are required to complete the study on a smartphone. We use this interface as it encourages participants to use our system as they would use predictive text in their everyday lives. Our mobile interface connects to a custom cloud-hosted back-end that uses a language model to provide predictive test suggestions.
We conduct a mixed between- and within-subjects study design. We assess the effects of writing without suggestions (control) versus with suggestions (treatments) in a between-subjects analysis. In the treatment condition, we evaluate the effects of various stereotypes in predictive text suggestions with a within-subjects analysis. We do not employ a fully between-subjects design as providing participants with only pro-stereotypical or anti-stereotypical suggestions may increase the chances that participants notice they are in a bias-centred study and change their writing behavior accordingly.
Our study procedure consists of up to two tutorial tasks, seven writing tasks, an attention check, a break and a final survey ordered as shown in Figure 2. After the study is completed, participants in the “with suggestions” (treatment) conditions receive a debrief explaining how the predictive text model was controlled in a way that influenced the character attributes the model suggested (e.g., suggesting that the doctor character in the story is a woman), leading to suggestions that may reinforce harmful stereotypes (see Figure 23).
Figure 2. Our study structure consists of a tutorial, seven writing tasks (with an additional attention check task), a short break, and a final survey. See Appendix D for a full set of interface screenshots.
[…]
@@ 4.1. Procedure
After finishing the tutorials, participants are asked to complete seven writing tasks (see 3(a)) the participant is given the starting words of a story and are asked to complete it. Participants are required to write at least 100 characters before they are able to move to the next scenario. We record an interaction trace of user behavior throughout each writing task. This includes every suggestion that is accepted or rejected by the user, every word they type or delete, and the amount of time taken on each of these actions. For our purposes, a writing action ends at a space character.
[…]
User interface. Please complete the following story: “I went to the doctor yesterday. The doctor said that”
@@
prefix I went to …. The doctor said that suggest [she, I] type “he” suggest [would, needed, prescribed] type “cannot” suggest [diagnose, determine, give] pick “diagnose” suggest [my, me, without] pick “me”
@@
I went to the doctor yesterday. The doctor saidgthathegcannotdiagnose me
[…]
@@
We employ two strategies to encourage participants strongly engage with the system and the writing task. First, we explain in the task instructions that participants’ usage of the predictive text is being monitored throughout the study and that their compensation may be affected if they exclusively and very quickly accept suggestions (in the end, all participants were compensated at the full rate). After the first scenario, if any, in which a participant writes more than $90\%$ of the words in the story via predictive suggestions, we include a warning screen reminding them not to over-use the suggestions. $28.5\%$ of participants in the suggestions condition were shown this warning. Being shown the warning did not affect participants’ compensation, ability to complete the study, or their inclusion in the analysis.
@@
Because our predictive text suggestions will attempt to “nudge” participants towards pro- or anti-stereotypical completions, we also collect a proxy measure of participants’ underlying beliefs (see Figure 21); see H3. As discussed in subsection 4.2, our study’s writing scenarios generally center an association between gender and another stereotype-relevant trait. These traits come from Koch et al. (2016)’s ABC model (building on Fiske et al. (2002)’s stereotype content model) which consists of paired traits regarding a group’s agency, beliefs, and communion. To measure the participant’s beliefs about these stereotypes, we ask one question about warmth (representing “communion”), competence (representing “agency”), and one about conservativeness (the only “belief” represented in our scenarios). For the conservativeness question, we use the proxy of “community-oriented” vs “individualistic” which aligns well with our liberal vs conservative writing scenario which focuses on affordable housing development. Similar to Cao et al. (2022b), we ask participants to mark on a 0-100 scale the extent to which different demographic groups are associated with warmth, competence, and conservativeness.
@@
4.2. Writing Scenarios
We present participants with seven writing scenarios to complete that involve various traits of interest (see Table 1). We carefully designed all scenarios (but one) to contain one gender and one Agency-Belief-Communion (Koch et al., 2016) axis. In some cases the gender is specified in the story prefix (e.g., the Student scenario) and in others, it is left to be determined by the person writing the story (e.g., the Doctor scenario). Similarly, in some cases, the ABC axis is specified in the story prefix (Detective) and in others it is left to be determined (Doctor). Importantly, and as discussed later, there is no guarantee that the person writing the story will always specify any of these in their story.
To dig in in more detail, in the Teacher scenario, participants start with the story prefix “When I was in school {Mr. / Mrs.} Brown was” where the teacher’s title is specified based on the condition. In this scenario, we also consider whether the teacher is likeable vs repellent (a “Communion” trait). On the other hand, in the President scenario, all stories begin with “In the first 100 days, the new president was determined to focus”. Here, we are concerned with the president’s gender and whether they are benevolent vs threatening (also “Communion” trait). While in the Teacher scenario, one axis (gender) was specified in the initial starting phrase of the story (and the likeability axes is possibly later specified by the participant), in the President scenario, both axes are left up to the participant.
The Wedding scenario is an exception where the two axes of interest are the gender of both of the characters who are getting married, and whether this is a straight or queer marriage. The seven scenarios are chosen to cover a wide variety of ABC traits and potential gender biases. The story prefixes are chosen to minimize the chance that a participant will immediately recognize the study’s focus on gender stereotypes. For example, if we marked characters as a “{male/female} doctor”, then participants may notice that the study is concerned with gender biases and adjust their writing accordingly.
Axes specified Axes suggested
Scenario Story prefix in story prefix by the model
@@ 4.2. Writing Scenarios
Teacher When I was in school, {Mr. / Mrs.} Brown was teacher gender teacher likeablity
@@ 4.3. Participants
$42\%$ of participants self-identified as women, $56\%$ as men, $1\%$ as non-binary/non-conforming, with $1\%$ of participants opting not to respond. $37\%$ of participants were between the ages of $18$-$25$, $43\%$ between $26$-$40$, $19\%$ between $41$-$60$, and $1\%$ over the age of $60$. $32\%$ of participants self-reported as having “primary fluency / bilingual proficiency” in English, $17\%$ as having “full professional proficiency”, $16\%$ as having “professional working proficiency”, $18\%$ as having “limited working proficiency”, and $16\%$ as having “elementary proficiency”.
@@ 5. Methods
Our study focuses on the effects of biases in an underlying predictive text model on participants’ behavior. In the study, participants write stories covering $7$ scenarios. In each scenario, participants are provided with an opening phrase and asked to continue the story. The underlying predictive text model (if any) can be biased in multiple ways, and we study the effects of that bias (if any) on the user-generated story.
@@ 5.1. Generating Predictive Text Suggestions
We generate our predictive text suggestions using Llama 2 7B (Touvron et al., 2023). Our model selection was based on a trade-off in ease and robustness of steering vs model size as we needed a model that would consistently suggest biased attributes as required but was also not so large as cause latency issues when making many word-level predictions. While Llama 2 7B may not be used in consumer predictive text systems, major companies have begun using transformer models for predictive text. [^5] https://www.apple.com/newsroom/2023/06/ios-17-makes-iphone-more-personal-and-intuitive/
@@
An example model prompt used in the study is shown in Figure 4. The majority of the system prompt is shared between scenarios and conditions. It explains the task of next word predictionand then, depending on the scenario and the condition, describes any aspects of the story that we are trying to control (e.g., explains that the model should suggest that a character is a woman). We then include two to three in-context examples showing how to continue a story with the desired characteristics. These sample continuations are generated in part with inspiration from gpt-3.5-turbo to help select diverse completions. Finally, we include the current state of the story as it is being written. [^6] The full set of prompts (with in-context examples) used in the study is available at [github link removed for anonymization].
We generate the top three predictive text suggestions using a simple decoding method. We start by generating the top three tokens. Since these tokens may not end on a meaningful word boundary, we continue greedily until each of the top three suggestions contains a completed word, digit, or punctuation mark. In the simplest case, this means we generate until we see a space. We also check for completed words containing apostrophes or hyphens (e.g., we should continue generating at “doctor”’ until we reach “doctor’s”). While these continuations may affect the probability of the full generation, we represent the probability of each suggestion based only on the probability of the first generated token as a small number of additional, greedily decoded tokens likely will not have a large impact on these probabilities.
5.2. Identification of Pertinent Story Elements
To identify whether stories’ characters have a particular gender or one of the ABC traits that is relevant to a scenario, we annotate the produced stories using the large language model (LLM) Llama3 70B(AI@Meta, 2024).
For example, in the co-written story “I went to the doctor yesterday. The doctor said that she would run additional tests to confirm the unpleasant results from the insulin levels to be true,” we want to know if the doctor is described using fem-coded language and if the doctor is confident or unconfident. We formulate this annotation as a Natural Language Inference task (Williams et al., 2018) where we provide the model with the story as a “premise” as well as a hypothesis such as “In the story, the doctor is a woman or a person who uses she/her pronouns or a traditionally feminine name,” (see Table 18 for the hypotheses used for every scenario). In this case, we expect the model to annotate this story as having a fem-coded doctor character. [^7] Throughout this paper, we use the terms masc-coded and fem-coded to represent characters that have lexical characteristics such as using certain pronouns. These characteristics are related but not equivalent to gender itself. We collect similar annotations at the word level, providing the model with the story up until a specific word (that may have actually been included in the story or suggested by the model and rejected) and evaluating the same hypotheses. Here, we expect to see the model’s probability of the doctor being fem-coded to increase significantly on the word “she”.
@@ 5.2. Identification of Pertinent Story Elements
[…]
Example predictive text prompt. The system prompt is “You are next word prediction model that helps people finish their current story. Please continue the story indefinitely as the user will decide when to stop, but feel encouraged to end individual sentences as needed. Only respond with the continuation with no extra text. The continuation should be from the same point of view as the original story. This story is about Detective Sarah Robinson who betrays her partner, Detective John Wilson. Again, remember to respond only with the continuation of the user’s story.” We then continue with in context examples of completing such a detective story and then provide the current state of the participant’s story.

### v2

@@ 3. Research Question and Hypotheses
We study the effect of biases in a predictive text system on co-writing creative stories. Participants in our study are assigned to either a control condition, in which they do not receive any text predictions, or the treatment condition, in which they do. In the treatment condition, as in standard phone keyboard interfaces, the participant is provided (up to) three predicted “next words” that they can select rather than typing on their own. The treatment condition can be further split based on the content of the model suggestions. Broadly, we have pro-stereotypical conditions where the model that provides word suggestions is configured to do so in a way that conforms to known social stereotypes and anti-stereotypical conditions where here the model is configured to provide suggestions that challenge social stereotypes. All stereotypes (pro- and anti-) are restricted to gender- and sexuality-based stereotypes.
@@
For example, in Figure 1, the model may suggest a president character should be described using masc-coded language (e.g., using he/him pronouns or having a traditionally masculine name; pro-stereotypical) or fem-coded language (e.g., using she/her pronouns or having a traditionally feminine name; anti-stereotypical). Beyond gender alone, the predictive text system may also suggest a number of gender-associated traits, for example that a fem-coded character is “benevolent” (pro-stereotypical as per cao-etal-2022-theory) or that she is “threatening” (anti-stereotypical as per cao-etal-2022-theory).
@@
Our analysis is primarily concerned with users’ decisions to accept or reject suggestions from a predictive text system (H2) and how these decisions lead to overall stories that are qualitatively similar or different from stories written without suggestions (H1). Measures of the acceptance of individual word-level suggestions capture different effects than measures of the degree of the use of stereotypes in the completed stories. The former provide measures of reliance. However, it is possible that simply observing the suggestions—without actually selecting them—influences what people write. Our story-level measures allow us to observe such influences at a holistic level.
The question of precisely what constitutes a “fair” outcome is essentially contested across multiple fields, including algorithmic fairness, philosophy, AI safety, and HCI (e.g., Narayanan, 2018; Mehrabi et al., 2021; Barocas et al., 2019; Chien et al., 2024; Holstein et al., 2019; Lee and Baykal, 2017). Many mathematical approaches cast “fairness” as some measure of disparity of outcomes across groups, with that precise measure also being essentially contested (Narayanan, 2018). Even the question of what level of disparities are acceptable is disputed: should no disparity be allowed, should disparities be allowed up to some real-world statistic (e.g. labor statistics from the country of model deployment), or something else? The “correct” definition of fairness certainly depends on both the goals of the designer and developer, as well as the actual context in which an AI-based system will be deployed.
We have hypothesized that providing anti-stereotypical suggestions will lead participants to write stories that are more anti-stereotypical than human-only stories, while providing pro-stereotypical suggestions may not result in a significant difference from human-only stories. We center parts of our analysis on whether the potential increase of anti-stereotypical stories from solely anti-stereotypical suggestions is enough to result in a “fair” distribution of stories–namely a distribution exhibiting demographic parity. However, this is not to say that demographic parity is the only reasonable fairness definition to apply in this setting, and we leave the governance question of what distribution of suggestions or final stories is “fair”—whether based on parity, real-world statistics, or other criteria—to future work.
@@
In contrast to studies of bias in language models that are either intrinsic or extrinsic to the model itself, these two hypotheses are concerned with how model biases affect co-writing with a human. H2 focuses on individual micro decisions about when participants accept model suggested words or reject them and write new words, [^4] Our hypotheses focus on suggested content being pro-stereotypical or anti-stereotypical as the difference maker that determines whether participants will accept or reject these suggestions. However, as we discuss in subsection B.6, another possible cause of differences in token-level acceptance of suggestions in our study is human preferences towards text with uniform information density (Frank and Jaeger, 2008; Jaeger and Levy, 2006; meister-etal-2021-revisiting). While we cannot rule out this potential confounder entirely, we discuss in subsection B.6 how participants often select markers of character gender sufficiently early in the co-writing process that our results cannot be explained by differences information density alone. and H1 focuses on the impact of those decisions to written stories more broadly.
@@
Beyond the main analyses introduced above, which we conduct in the main body of this paper (section 6), we conduct a few additional analyses in the appendices. In these additional analyses, we observe that: (1) the presence of suggestions did not affect the overall story lengths (subsection B.2); (2) participants’ (binary) gender identity did not significantly affect their acceptance of gendered suggestions (subsection B.3); (3) human biases correlated with each other, for example, with groups being seen as “warm” also being seen as “competent” (subsection B.4); and (4) writing with predictive text did not significantly affect gender gaps in toxicity but led to significant gender gaps in sentiment and character agency in some writing scenarios, with fem-coded characters sometimes being portrayed more positively, yet with less agency. (subsection B.5).
@@ 4. Study Design
The study is conducted using a custom-built mobile web interface (3(a)) mimicking a smartphone keyboard with predictive text, and participants are required to complete the study on a smartphone and were not allowed to use their device keyboard. We use this interface as it encourages participants to use our system as they would use predictive text in their everyday lives. Our mobile interface connects to a custom cloud-hosted back-end that uses a language model to provide predictive text suggestions.
We use a mixed between- and within-subjects study design. We assess the effects of writing without suggestions (control) versus with suggestions (treatments) in a between-subjects analysis. In the treatment condition, the stereotypes present in the predictive text suggestions vary within-subjects and are randomized for each writing task. We do not employ a fully between-subjects design as providing only pro-stereotypical or anti-stereotypical suggestions may increase the chances that participants notice they are in a bias-centered study and change their writing behavior accordingly.
Our study procedure consists of up to two tutorial tasks, seven writing tasks, an attention check, a break and a final survey ordered as shown in Figure 2. After the study is completed, participants who received predictive text suggestions were shown a debrief explaining how the predictive text model was controlled in a way that influenced the character attributes the model suggested (e.g., suggesting that the doctor character in the story is a woman), leading to suggestions that may reinforce harmful stereotypes (See Figure 26).
Figure 2. Our study structure consists of a tutorial, seven writing tasks (with an additional attention check task), a short break, and a final survey. See Appendix E for a full set of interface screenshots.
@@ 4.1. Procedure
After finishing the tutorials, participants are asked to complete seven writing tasks (See 3(a)) in which the participant is given the opening words of a story and are asked to complete it. Participants are required to write at least 100 characters before they are able to move to the next scenario. We record an interaction trace of participant behavior throughout each writing task. This includes every suggestion that is accepted or rejected by the participant, every word they type or delete, and the amount of time taken on each of these actions. For our purposes, a writing action ends at a space character.
@@
prefix I went to …. The doctor said that
suggest [she, I]
type “he”
suggest [would, needed, prescribed]
type “cannot”
suggest [diagnose, determine, give]
pick “diagnose”
suggest [my, me, without]
pick “me”
@@
I went to the doctor yesterday. The doctor
saidgthathegcannotdiagnose me
@@
(a) Screenshot of the writing interface showing the prompt “I went to the doctor yesterday. The doctor said that.” (b) Stylized interaction trace illustrating a participant typing words and selecting predictive text suggestions while completing the sentence, resulting in “The doctor said that he cannot diagnose me.” The trace illustrates alternating model suggestions and participant selections.
We employ two strategies to encourage participants strongly engage with the system and the writing task. First, we explain in the task instructions that participants’ usage of predictive text is being monitored throughout the study and that their compensation may be affected if they exclusively and very quickly accept suggestions (in the end, all participants were compensated at the full rate). After the first scenario, if any, in which a participant writes more than $90\%$ of the words in the story via predictive suggestions, we include a warning screen reminding them not to overuse the suggestions. $28.5\%$ of participants in the suggestions condition were shown this warning. Being shown the warning did not affect participants’ compensation, ability to complete the study, or their inclusion in the analysis.
@@
Because our predictive text suggestions will attempt to “nudge” participants towards pro- or anti-stereotypical completions, we also collect a proxy measure of participants’ underlying beliefs (See Figure 24). As discussed in subsection 4.2, our study’s writing scenarios generally center an association between gender and another stereotype-relevant trait. These traits come from Koch et al. (2016)’s ABC model (building on Fiske et al. (2002)’s stereotype content model) which consists of paired traits regarding a group’s agency, beliefs, and communion. To measure the participant’s beliefs about these stereotypes, we ask one question about warmth (representing “communion”), competence (representing “agency”), and one about conservativeness (the only “belief” represented in our writing scenarios). For the conservativeness question, we use the proxy of “community-oriented” vs “individualistic” which aligns well with our liberal vs conservative writing scenario which focuses on affordable housing development. Similar to cao-etal-2022-theory, we ask participants to mark on a 0-100 scale the extent to which different demographic groups are associated with warmth, competence, and conservativeness.
@@
Scenario Story prefix Axes specified in story prefix Axes sugggested by the model
@@
Teacher When I was in school, {Mr. / Mrs.} Brown was teacher gender teacher likablity
@@
4.2. Writing Scenarios
We present participants with seven writing scenarios to complete that involve various traits of interest (See Table 1). All scenarios (except one) were designed to include one gender and one Agency-Belief-Communion (Koch et al., 2016) axis (distributed as two Agency, three Belief, and one Communion scenario), with some scenarios specifying one of these attributes in the story prefix (e.g., the Student scenario) and others leaving them open to the writer (e.g., the Doctor scenario). These differences allow us to probe how participants assign traits in contexts where gender or ABC traits are either given or unspecified. For example, in the Detective scenario, we test whether a detective is more likely to be gendered masculine or feminine based on the detective being explicitly marked as trustworthy or untrustworthy. In the Doctor, neither gender nor the ABC trait (confidence) is specified in the story prefix, so we can see which which combinations of traits are chosen jointly by the participants. The Wedding scenario does not involve ABC traits. Instead, the two axes of interest are the gender of both of the characters who are getting married, and whether this is a straight or queer marriage.
These scenarios also differ in the real-world base-rates of attribute combinations—for example, the proportion of female presidents is lower than that of female doctors—though for ABC traits, no meaningful population statistics can reasonably be obtained (e.g., the proportion of female doctors who are confident vs. unconfident). We do not strictly control these rates. Instead, these variations provide a range of contexts that may influence the strength of participants’ stereotypical beliefs. Importantly, and as discussed later, there is no guarantee that the person writing the story will always specify any of these attributes (gender or ABC trait) in their stories.
To dig in in more detail, in the Teacher scenario, participants start with the story prefix “When I was in school {Mr. / Mrs.} Brown was” where the teacher’s title is specified based on the condition. In this scenario, we also consider whether the teacher is likable vs repellent (a “Communion” trait). On the other hand, in the President scenario, all stories begin with “In the first 100 days, the new president was determined to focus”. Here, we are concerned with the president’s gender and whether they are benevolent vs threatening (also “Communion” trait). While in the Teacher scenario, one axis (gender) was specified in the initial starting phrase of the story (and the likability axes is possibly later specified by the participant), in the President scenario, both axes are left up to the participant.
Overall, the seven scenarios are chosen to cover a wide variety of ABC traits and potential gender biases. The story prefixes are chosen to minimize the chance that a participant will immediately recognize the study’s focus on gender stereotypes. For example, if we marked characters as a “{male/female} doctor”, then participants may notice that the study is concerned with gender biases and adjust their writing accordingly.
@@ 4.3. Participants
$42\%$ of participants self-identified as women, $56\%$ as men, $1\%$ as non-binary/non-conforming, with $1\%$ of participants opting not to respond. $37\%$ of participants were between the ages of 1818-$25$, $43\%$ between 2626-$40$, $19\%$ between 4141-$60$, and $1\%$ over the age of $60$. $32\%$ of participants self-reported as having “primary fluency / bilingual proficiency” in English, $17\%$ as having “full professional proficiency”, $16\%$ as having “professional working proficiency”, $18\%$ as having “limited working proficiency”, and $16\%$ as having “elementary proficiency”. Participation was not restricted by country of origin to ease the recruitment of participants with a variety of English proficiency levels. A breakdown of participant nationality can be seen in Figure 20.
@@ 5. Methods
Our study focuses on the effects of biases in an underlying predictive text model on participants’ behavior. In the study, participants write stories covering seven scenarios. In each scenario, participants are provided with an opening phrase and asked to continue the story. The underlying predictive text model (if any) can be biased in multiple ways, and we study the effects of that bias (if any) on the user-generated story.
@@ 5.1. Generating Predictive Text Suggestions
We generate our predictive text suggestions using Llama 2-Chat 7B (Touvron et al., 2023). Our model selection was based on a trade-off in ease and robustness of steering vs model size as we needed a model that would consistently suggest biased attributes as required but was also not so large as to cause latency issues when making many word-level predictions. While Llama 2 7B may not be used in consumer predictive text systems, major companies have begun using transformer models for predictive text. [^6] https://www.apple.com/newsroom/2023/06/ios-17-makes-iphone-more-personal-and-intuitive/
@@
An example model prompt used in the study is shown in Figure 4. The majority of the system prompt is shared across scenarios and conditions. It explains the next-word prediction task and then, depending on the scenario and condition, describes specific aspects of the story we aim to control (e.g., that the model should suggest a character is a woman). We then include two to three in-context examples showing how to continue a story with the desired characteristics. These sample continuations are generated in part with inspiration from gpt-3.5-turbo to help select diverse completions. Finally, we include the current state of the story as it is being written. [^7] The full set of prompts (with in-context examples) used in the study is available at https://github.com/ctbaumler/predictive_text.
@@
Example predictive text prompt showing the system instruction to continue a story indefinitely, followed by in-context examples of completing a detective story about Detective Sarah Robinson and Detective John Wilson, and the current state of the participant’s story.
We generate the top three predictive text suggestions using a simple decoding method. We start by taking the top three tokens according to their raw output logits. Since these tokens may not end on a meaningful word boundary, we continue greedily until each of the top three suggestions contains a completed word, number, or punctuation mark. In the simplest case, this means we generate until we see a space. We also check for completed words containing apostrophes or hyphens (e.g., we should continue generating at “doctor”’ until we reach “doctor’s”). While these continuations may affect the probability of the full sequence, we approximate the probability of each suggestion using only the probability of the first generated token. Because only a small number of additional tokens are generated greedily, their contribution is unlikely to substantially change the relative ranking of the suggestions.
For more details about how predictive text suggestions were generated, see Appendix D.
5.2. Identification and Measurement of Pertinent Story Elements
To identify whether stories’ characters have a particular gender or one of the ABC traits that is relevant to a scenario, we annotate the produced stories using Llama3 70B(AI@Meta, 2024).
For example, in the co-written story “I went to the doctor yesterday. The doctor said that she would run additional tests to confirm the unpleasant results from the insulin levels to be true,” we want to know if the doctor is described using fem-coded language and if the doctor is described as confident or unconfident. We formulate this annotation as a Natural Language Inference task (williams-etal-2018-broad) in which we provide the model with the story as a “premise” as well as a hypothesis such as “In the story, the doctor is a woman or a person who uses she/her pronouns or a traditionally feminine name,” (See Table 18 for the hypotheses used for every scenario) and collect the probability of the hypothesis being true via the resulting token probabilities. We expect the model to annotate above story as having a fem-coded doctor character. [^8] Throughout this paper, we use the terms masc-coded and fem-coded to represent characters that have lexical characteristics such as using certain pronouns. These characteristics are related but not equivalent to gender itself. We collect similar annotations at the word level, providing the model with the story up until a specific word (that may have actually been included in the story or suggested by the model and rejected) and evaluating the same hypotheses. Here, we expect to see the model’s probability of the doctor being fem-coded to increase significantly on the word “she”.